### v1

@@ Abstract
Dirichlet process mixture models (DPMM) provide a principled framework for density estimation and data clustering, offering a robust alternative to parametric mixture models by estimating cluster number $K$ from an infinite choice of partitions induced by a Dirichlet Process $DP(\alpha,G_{0})$ prior on the data. Posterior consistency is guaranteed by an adaptive estimation of the DP concentration parameter $\alpha$, and the estimated $K$ jointly depends on $\alpha$ as well as the variance hyper-parameter of DP base distribution $G_{0}$. Although Markov Chain Monte Carlo (MCMC) methods have bridged the gap between theory and application for such models, they scale poorly to high-dimensional data (e.g. with several hundred or even thousands of features) with limited sample size and suffer from slow and difficult convergence. Variational Inference (VI) exist as an alternative for faster convergence, but lack implementation with increasing model fidelity. Collapsed VI is implemented to integrate out hierarchical parameters beyond mean field VI’s independence assumptions, however, the resulting complexity imposes an implementational barrier to a fully adaptive DPMM inference for high dimensional data. We propose a novel method that performs adaptive clustering with DPMM using collapsed VI, while incorporating weakly-informative priors for $\alpha$ and $G_{0}$. We illustrate the importance of $G_{0}$ covariance structure and prior choice by considering different parameterisations of the data covariance matrix. On high-dimensional Gaussian simulations, our model demonstrates substantially faster convergence than a state-of-the-art MCMC splice sampler. We further evaluate performances on Negative Binomial simulations and conduct sensitivity analyses to assess robustness on realistic data conditions. Application to a publicly available leukemia transcriptomic data set comprising 72 samples and 2,194 gene expression successfully recovers every known sub-type, all while identifying additional gene expression-based sub-clusters with meaningful biological interpretation.
keywords:
Clustering, Bayesian Nonparametrics, Dirichlet process mixture model, Variational inference, Unstructured covariance, Concentration parameter, High-dimension, Gene expression data
issue: x
@@ 1 Introduction
Clustering is an exploratory data analysis technique, where the goal is to group the data based on measured or estimated inherent similarities (jain2010data). Since the early emergence of $K$-means in 1950s (steinhaus1956division; forgy1965cluster; mcqueen1967some; lloyd1982least), a large number of clustering methods have been deployed for diverse interdisciplinary applications that deal with high dimensional data sets (dinh2025data). These methods include density based approaches (ester1996density; campello2013density), graph based techniques (ertoz2002new; traag2019louvain) and probabilistic models that define clusters based on mixture distributions (bock1996probabilistic, bouveyron2007high). However, a prerequisite for standard clustering techniques and finite mixture models is the number of clusters $K$ (or analogous parameters for density based methods) present a priori (cai2021finite). Even dimensionality reduction techniques like PCA (pearson1901liii) or UMAP (mcinnes2018umap) fail to capture underlying structure for high-dimensional data and require strong domain-specific knowledge to select an appropriate $K$ (wani2025comprehensive). To circumvent this limitation, Bayesian non-parametric mixture models (NPMM) have emerged as an efficient alternative to perform adaptive clustering based on probability distributions without pre-specifying $K$ (orbanz2010bayesian; gershman2012tutorial).
@@
Introduction of Dirichlet process $DP(\alpha,G_{0})$ by ferguson1973bayesian enabled the development of DP mixture models (DPMM) as a class of Bayesian NPMM for density estimation, clustering as well as regression (muller2013bayesian). DPMM assume an infinite choice of unknown $K$ corresponding to partitions of the mixture components based on a DP prior (fruhwirth2019here), and corresponding posterior inference facilitates the estimate of $K$ (a non-empty partition corresponds to a cluster $k\in K$). Although DPMM have been scrutinized for inconsistent posterior estimation (miller2014inconsistency; miller2013simple), ascolani2023clustering have successfully demonstrated that consistency can be obtained by varying the concentration parameter $\alpha$, for instance, using a Gamma prior to implement a fully Bayesian scheme to update $\alpha$ (west1992hyperparameter).
@@
DPMM often results in analytically intractable likelihoods and posteriors. Markov Chain Monte Carlo (MCMC) methods are a popular choice to sample from such analytically intractable distributions for parameter inference (escobar1995bayesian; neal2000markov). However, these methods exhibit slow convergence rates with difficulty in assessment (gershman2012tutorial), especially for high dimensions with several hundreds or even thousands of features per sample and limited sample size. An alternative is to use variational inference (VI) that approximates the true posterior $p(\cdot|data)$ with a tractable density function $q(\cdot)$ from a (known) family of probability distributions while deterministically optimising an information-theoretic criterion like KL divergence (blei2017variational). Rooted in early estimation methods for graphical models (jordan1999introduction), VI has gained prominence for its computational efficiency and scalability, particularly in modern machine learning applications involving large datasets (bernardi2024variational; loya2025scalable).
@@
A general scheme of DPMM implementation would involve a stick-breaking representation of the DP prior (sethuraman1994constructive) with hyper-priors defined for $\alpha$ and base distribution $G_{0}$. To estimate model parameters, mean-field VI is the most widely adopted approach that assumes a factorized variational distribution and iteratively updates the parameters using conjugate priors (blei2006variational; bishop2006pattern). Although the independent factorization among variational parameters induces analytical tractability, latent allocation variables in DPMM are inherently dependent through their shared mixture proportions across samples. This fails to capture the relationships between hierarchical parameters and results in inconsistency between cluster numbers estimated from $\alpha$ and the latent allocation variables respectively (details in Supplementary material). kurihara2007collapsed have provided a collapsed VI framework by integrating out (collapsing) the mixture proportions, thus defining a prior distribution for the latent allocation vectors that depends only on $\alpha$ (1). This integration, however, constructs a non-standard prior with no standard conjugate hyper-prior for $\alpha$. Hence, the collapsed VI approach lacks a current working implementation with an adaptive inference on $\alpha$.
@@
Apart from $\alpha$ that has a directly proportional effect on the inferred number of clusters (li2019tutorial), the model variance also plays a significant role in estimating $K$ (hejblum2019sequential). Incorporating $G_{0}$ variance as an unknown variable becomes important for coherent estimation, which is often ignored while implementing DPMM (including blei2006variational). The cluster variance particularly impacts high-dimensional data, which is often characterized by reduced signal-to-noise ratio, high variability due to limited sample size (johnstone2009statistical) and reduced separation between data points with increasing dimensions (clarke2008properties). The commonly used Inverse-Wishart $IW$ prior for covariance matrix receives criticism for a rigid structure, strongly coupled relations between variance and covariance (o2004kendall; tokuda2025visualizing) and high sensitivity to hyper-parameter choice (hennig2015handbook). Factorized covariance representations (like Cholesky decomposed factors) exhibit similar variance-covariance dependency as the resulting matrix have linear combination of constituent factor elements that are common in variance as well as covariance terms. Approaches like block diagonal covariance matrices, latent factor models (chandra2023escaping) and element-wise distributions (jing2024variance) focus on introducing conditional independence within the covariance matrices. A fundamental issue with Gaussian mixture models or GMM (both finite and infinite) is their assumption of Gaussian distributed clusters. This assumption is frequently violated in real world data sets, where cluster distributions may exhibit skewness, heavy tails, multimodality, or other deviations from normality (hejblum2019sequential; zhang2022bayesian). Paired with the ’curse of dimensionality’, data sets like gene expression data often show less separated or overlapped clusters (yu2017clustering) and using GMMs lead to misspecified clustering (kasa2023avoiding). In biological applications, the cluster overlapping can also arise from factors like multimodality and high correlation (clarke2008properties). In order to retrieve clustering estimates from such data, it is important to have cluster-specific distributions that regulate within-cluster spread as well as inter-cluster separability for robust outcomes.
@@
In this paper, we propose a novel collapsed VI approach that explicitly incorporates the concentration parameter $\alpha$ as well as model covariance in a DPMM. Combining Euler-Maclaurin and Taylor series approximations, we obtain a closed form variational distribution for $\alpha$ with a conjugate Gamma hyper-prior. Following a selection between different covariance parameterisations, the analysis evaluates effect of hyper-prior choice through their corresponding performances on Gaussian simulated data. We establish the model with sparsity inducing hyper-prior (namely, Sparse DPMM) as the best choice (a desirable property for high-dimensional as well as highly-clustered data (jing2024variance)) and further evaluate the sensitivity based on the hyper-parameters with high dimensional ($d=1000$) Negative Binomial simulations. The optimal model is compared to an MCMC approach as implemented in the R package NPflow (hejblum2019sequential) in terms of convergence speed. We apply Sparse DPMM to cluster leukemia sub-types based on gene expression data set (armstrong2002mll; de2008clustering) with $2194$ genes and $72$ samples, and further evaluated the performance with existing clustering techniques. In addition, we provide a computational implementation through the R-package vimixr, available on CRAN.
The rest of the paper continues as follows. Section 2 develops the statistical methodology of our proposed model, focusing on formulating variational updates for different $G_{0}$ covariance hyper-priors and model implementation (Table 1). Section 3 highlights the theoretical merits and illustrates the inference results on simulated data for optimal model choice and sensitivity analysis with an emphasis on hyper-parameter tuning, while comparing computational performance with a splice sampling MCMC approach (hejblum2019sequential). It is followed by an application of Sparse DPMM on gene expression data to cluster acute lymphoblastic (ALL), mixed lineage (MLL) and acute myeloid (AML) leukemia sub-types (armstrong2002mll) and compare performance with $K$-means (mcqueen1967some), density-based DBSCAN (ester1996density) and HDBSCAN (campello2013density), graph-based shared nearest neighbourhood (ertoz2002new) and Leiden (kelly2023leiden), and model-based high dimensional clustering (bouveyron2007high) techniques. Section 4 concludes with discussions on the benefits and short comings of the proposed model, with possible extensions for further research and development.
@@ 2.1 Problem set-up
where $F(\mathbf{X})=\int_{\Theta}f_{\eta}(\mathbf{X})G(d\eta)$ (hejblum2019sequential). We assume a multivariate Gaussian distribution for $f_{\eta}(.)$ with mean $\mu$ and covariance matrix $\Sigma$ ($\eta=\{\mu,\Sigma\}$). $G$ is the unknown mixing distribution that characterizes the mixture components over the parameter space $\Theta$, thus inducing clustering among $X_{n}$. A DP prior on $G\sim DP(\alpha,G_{0})$ with concentration parameter $\alpha$ and base distribution $G_{0}$ (ferguson1973bayesian) gives us a non-parametric mixing distribution
@@
The mixing proportions $\pi_{k}$’s are drawn from a stick-breaking scheme (sethuraman1994constructive), which further introduce latent allocation variables $Z_{n}$ in the model that follows a categorical distribution with parameters $\{\pi_{k}\}$’s such that $p(Z_{n}=k)=\pi_{k}$ for every $n\in N$. The likelihood of $X_{n}$ hence follows
@@
(blei2006variational). Our objective is to estimate $Z_{n}$ as well as $\eta_{k}=\{\mu_{k},\Sigma_{k}\}$ to quantitatively evaluate the partitioning of the parameter space $\mathbf{\Theta}$.
@@ 2.2 Dirichlet process mixture model
This is an illustration of a Dirichlet process mixture model (DPMM), introduced by antoniak1974mixtures. The concentration parameter $\alpha$ has a direct impact on the posterior expectation of the number of non-empty clusters (teh2010dirichlet). escobar1995bayesian have proposed a data augmentation scheme that leverages a Gamma hyper-prior $\alpha\sim Gamma(a,b)$ yielding a posterior distribution that adjust to the actual number of clusters observed in the data with consistency (ascolani2023clustering). With a conjugate choice for $G_{0}$, the model provides an estimate for the unknown parameters of interest.
@@ 2.3 Choice of $G_{0}$
The underlying structure of $\eta=\{\mu,\Sigma\}$ governs the distributional choice for $G_{0}$. We use a multivariate Gaussian $MVN(\mathbf{0},\Sigma_{\mu})$ as the prior for $\mu_{k}$. To illustrate the effect of variance parameter on a clustering model, several parameterisation for $\Sigma$ are presented in Table 1.
@@
We introduce gradual complexity in terms of covariance matrix choices, starting from known or fixed variance ($M_{1},M_{2}$). For unknown variance, global refers to a common covariance matrix $\Sigma$ for all cluster choices ($M_{3},M_{4},M_{5}$), whereas cluster-specific defines each cluster with a unique covariance matrix $\Sigma_{k}$, along with the mean vectors ($M_{6},M_{7},M_{8}$). The choice of prior distributions maintain the conjugacy of the models, while considering hyper-parameters that establish weakly informative priors.
@@ 2.4 Posterior estimation using collapsed VI
We implement a collapsed variational inference (CVI) approach, where we integrate out (or collapse) the stick-breaking parameters $\{V_{k}\}$, following kurihara2007collapsed. This yields a distribution of the latent allocation vectors $\{z_{n}\}$ depending only on the DP concentration parameter $\alpha$:
@@
(blei2017variational; zhang2018advances). Although having theoretically “infinite“ prior choices for $k$, truncated variational distributions with an upper limit $k=K$ enable practical implementation (ishwaran2001gibbs).
For posterior estimation, we use exponential families of distributions for $q(\cdot)$ and apply coordinate ascent algorithm to update the variational hyper-parameters (bishop2006pattern). It ensures guarantee of convergence and provides a closed form update $q^{*}(\cdot)$ that maximizes the Evidence lower bound (ELBO) (or equivalently minimizing the information-theoretic Kullbach-Leibler (KL) divergence)
@@
where $\theta_{j}\in\mathbf{\Theta}=\{\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\}\}$ and $\theta^{-j}=\mathbf{\Theta}\backslash\theta_{j}$ represents the set of all except the $j^{th}$ parameter (blei2006variational; bishop2006pattern). Appropriate prior as well as variational distribution choices facilitate the update of $q^{*}(\cdot)$ hyper-parameters. These updates are used to perform posterior inference for the unknown $\{z_{n}\},\{\theta_{i}\}$, thus estimating data cluster as well as distributional partitions corresponding to the cluster allocations.

### v2

@@ Abstract
Dirichlet process mixture models (DPMM) provide a flexible framework for clustering complex biological data without pre-specifying the number of clusters. Although Markov Chain Monte Carlo (MCMC) methods have bridged the gap between theory and application for such models, they scale poorly to high-dimensional data (e.g. omics data) and suffer from slow and difficult convergence. Variational Inference (VI) represents a faster alternative, but lacks implementation without overly simplifying assumptions such as mean field independence and known covariance. We propose a novel methodology that performs scalable inference of DPMM for clustering using collapsed VI, ensuring coherent and data-driven estimation of the number of clusters in a fully Bayesian scheme. To accommodate high-dimensional data, we introduce cluster-specific adaptive covariance modeling with sparsity-inducing priors, improving flexibility while mitigating variance–covariance coupling and overparameterization. The resulting algorithm scales efficiently with dimension and sample size, substantially reducing computation time compared to state-of-the-art MCMC splice sampler. Realistic simulation studies under Gaussian and negative-binomial settings demonstrate accurate recovery of cluster structure and robust performance across varying signal-to-noise regimes, supported by sensitivity analyses. Application on a publicly available leukemia transcriptomic dataset comprising 72 samples and 2,194 gene expressions successfully recovers established subtypes while revealing biological plasticity of one mixed lineage leukemia outlier sample. The proposed framework provides a computationally efficient and theoretically coherent solution for high-dimensional Bayesian clustering with unknown concentration and adaptive covariance.
Keywords: Bayesian non-parametrics; Dirichlet process mixture model; Gene expression data; High-dimensional clustering; Variational inference.
@@ 1 Introduction
Clustering is an exploratory data analysis technique, where the goal is to group the data based on measured or estimated inherent similarities (Jain, 2010). Since the early emergence of $K$-means from 1956 onward (Steinhaus and others, 1956; MacQueen, 1967), clustering has become increasingly prevalent in the analysis of complex biological datasets to identify and characterize their underlying structure, and a large number of methods have been deployed to deal with clustering high dimensional datasets. These methods include density based approaches (Ester et al., 1996; Campello et al., 2013), graph based techniques (Ertoz et al., 2002; Traag et al., 2019) and probabilistic models that define clusters based on mixture distributions (Bock, 1996; Bouveyron et al., 2007). However, a prerequisite for standard clustering techniques and finite mixture models is the number of clusters $K$ (or analogous parameters for density as well as graph based methods) present a priori. Even dimensionality reduction techniques like PCA (Pearson, 1901) and UMAP (McInnes et al., 2018) often fail to capture underlying structure for high-dimensional data and require strong domain-specific knowledge to select an appropriate $K$ (Wani, 2025). To circumvent this limitation, Bayesian non-parametric mixture models (NPMM) have emerged as an efficient alternative to perform clustering based on probability distributions without pre-specifying $K$ (Gershman and Blei, 2012).
@@
Ferguson (1973) introduced Dirichlet process $DP(\alpha,G_{0})$ that led to the development of DP mixture models (DPMM) as a class of Bayesian NPMM for density estimation, clustering and regression (Müller and Mitra, 2013). DPMM use an infinite dimensional parameter space over the partition space of the observations thanks to a DP prior, yielding a posterior distribution with varying $K$. Although DPMM have been scrutinized for inconsistent posterior estimation of $K$ ($K$ being the number of non-empty clusters) by Miller and Harrison (2013), Ascolani et al. (2023) successfully demonstrated that consistency is guaranteed as long as the concentration parameter $\alpha$ is not fixed but data-driven (Escobar and West, 1995).
@@
DPMM often results in analytically intractable likelihoods and posteriors. Markov Chain Monte Carlo (MCMC) methods are a popular choice to sample from such distributions for parameter inference (Neal, 2000). However, these methods exhibit slow convergence rates with difficulty in assessment (Gershman and Blei, 2012), especially for high dimensions with several hundreds or even thousands of features per sample and limited sample size. An alternative is to use variational inference (VI) that approximates the true posterior $p(\cdot|data)$ with a tractable density function $q(\cdot)$ from a (known) family of probability distributions, while optimising an information-theoretic criterion like KL divergence (Blei et al., 2017). Rooted in early estimation methods for graphical models (Jordan et al., 1999), VI has gained prominence for its computational efficiency and scalability, particularly in modern machine learning applications involving large datasets (Bernardi et al., 2024; Loya et al., 2025).
@@
A general scheme of DPMM implementation would involve a stick-breaking representation of the DP prior (Sethuraman, 1994) with hyper-priors defined for $\alpha$ and base distribution $G_{0}$. For estimation, mean-field VI is a widely adopted approach that assumes factorized variational distribution and iteratively updates the parameters using conjugate priors (Blei and Jordan, 2006; Bishop and Nasrabadi, 2006). Although the independent factorization induces analytical tractability, latent allocation variables in DPMM are inherently dependent through their shared mixture proportions across samples. This fails to capture the relationships between hierarchical parameters and results in inconsistency between cluster numbers estimated from $\alpha$ and the latent allocation variables respectively (details in Supplementary material). Kurihara et al. (2007) provided a collapsed VI framework by integrating out (collapsing) the stick-breaking proportions, thus defining a prior distribution for the latent allocation vectors that depends only on $\alpha$ (1). This integration, however, constructs a non-standard prior with no standard conjugate hyper-prior for $\alpha$. Hence, the collapsed VI approach lacks a current working implementation with an adaptive inference on $\alpha$.
@@
Apart from $\alpha$ that has a directly proportional effect on the inferred number of clusters (Teh and others, 2010), the model variance also plays a significant role in estimating $K$ (Frühwirth-Schnatter and Pyne, 2010; Hejblum et al., 2019). Incorporating $G_{0}$ variance as an unknown variable becomes important for coherent estimation, which is often ignored while implementing DPMM (including Blei and Jordan (2006)). The cluster variance particularly impacts high-dimensional data, which is often characterized by reduced signal-to-noise ratio, high variability due to limited sample size, and reduced separation between data points with increasing dimensions (Clarke et al., 2008). Commonly used Inverse-Wishart $IW$ prior receives criticisms for a rigid and too informative structure and strongly coupled relations between variance and covariance (Tokuda et al., 2025). Factorized covariance representations (like Cholesky decomposed factors) exhibit similar variance-covariance dependency as the resulting matrix share factor elements by linear combinations. Alternative approaches are block diagonal covariance matrices, latent factor models and element-wise distributions (Jing et al., 2024) that focus on introducing conditional independence within the covariance matrices. Although widely used, a fundamental issue with Gaussian mixture models (GMM) is that cluster distributions may exhibit significant deviations from normality (Hejblum et al., 2019), particularly in real-world data. Paired with the ’curse of dimensionality’, datasets like transcriptomics data often show less separated or overlapped clusters (Yu et al., 2017) and using GMM lead to misspecified clustering. In biological applications, the cluster overlapping can also arise from factors like multimodality and high correlation (Clarke et al., 2008). In order to retrieve clustering estimates from such data, it is important to have cluster-specific distributions that regulate within-cluster spread as well as inter-cluster separability for robust outcomes.
@@
In this paper, we propose a novel collapsed VI approach that adaptively incorporates the concentration parameter $\alpha$ as well as model covariance in a DPMM. Combining Euler-Maclaurin and Taylor series approximations, we obtain a closed form variational distribution for $\alpha$ with a conjugate Gamma hyper-prior. Following a selection between different covariance parameterizations (Table 1), we formulate corresponding variational updates and evaluate the effect of hyper-prior choice through their performances on Gaussian simulated data. We establish the model with sparsity inducing hyper-prior (namely, Sparse DPMM) as the best choice (a desirable property for high-dimensional as well as highly-clustered data (Jing et al., 2024)) and further evaluate hyper-parameter sensitivity with high dimensional Negative Binomial simulations. The optimal model is compared to an MCMC approach as implemented in the R package NPflow (Hejblum et al., 2019) in terms of convergence speed. We apply Sparse DPMM to cluster acute lymphoblastic (ALL), mixed lineage (MLL) and acute myeloid (AML) leukemia sub-types based on gene expression dataset (Armstrong et al., 2002; De Souto et al., 2008) with $2194$ genes and $72$ samples, and benchmark the performance with existing $K$-means (MacQueen, 1967), density-based DBSCAN (Ester et al., 1996) and HDBSCAN (Campello et al., 2013), graph-based shared nearest neighbourhood (Ertoz et al., 2002) and Leiden (Traag et al., 2019), and model-based high dimensional clustering (Bouveyron et al., 2007) techniques. Sparse DPMM shows better recovery of known clusters than the conventional methods, while identifying an additional sub-cluster that corresponds to the plastic nature of MLL sub-type. Finally, we conclude with a general discussion of the proposed model, with possible directions for future research. We also provide a computational implementation through the R-package vimixr, available on CRAN.
@@ 2.1 Problem set-up
where $F(\mathbf{X})=\int_{\Theta}f_{\eta}(\mathbf{X})G(d\eta)$ (Hejblum et al., 2019). We assume a multivariate Gaussian distribution for $f_{\eta}(.)$ with mean $\mu$ and covariance matrix $\Sigma$ ($\eta=\{\mu,\Sigma\}$). $G$ is the unknown mixing distribution that characterizes the mixture components over the parameter space $\Theta$, thus inducing clustering among $X_{n}$. A DP prior on $G\sim DP(\alpha,G_{0})$ with concentration parameter $\alpha$ and base distribution $G_{0}$ (Ferguson, 1973) gives us a non-parametric mixing distribution
@@
The mixing proportions $\pi_{k}$’s are drawn from a stick-breaking scheme (Sethuraman, 1994), which further introduce latent allocation variables $Z_{n}$ in the model that follows a categorical distribution with parameters $\{\pi_{k}\}$’s such that $p(Z_{n}=k)=\pi_{k}$ for every $n\in N$. The likelihood of $X_{n}$ hence follows
@@
(Blei and Jordan, 2006). Our objective is to estimate $Z_{n}$ as well as $\eta_{k}=\{\mu_{k},\Sigma_{k}\}$ to quantitatively evaluate the partitioning of the parameter space $\mathbf{\Theta}$.
@@ 2.2 Dirichlet process mixture model
This is an illustration of a Dirichlet process mixture model (DPMM), introduced by Antoniak (1974). The concentration parameter $\alpha$ has a direct impact on the posterior expectation of the number of non-empty clusters (Teh and others, 2010). Escobar and West (1995) have proposed a data augmentation scheme that leverages a Gamma hyper-prior $\alpha\sim Gamma(a,b)$ yielding a posterior distribution that adjust to the actual number of clusters observed in the data with consistency (Ascolani et al., 2023). With a conjugate choice for $G_{0}$, the model provides an estimate for the unknown parameters of interest.
@@ 2.3 Choice of $G_{0}$
The underlying structure of $\eta=\{\mu,\Sigma\}$ governs the distributional choice for $G_{0}$. We use a multivariate Gaussian $MVN(\mathbf{0},\Sigma_{\mu})$ as the prior for $\mu_{k}$. To illustrate the effect of variance parameter on a clustering model, several parameterization for $\Sigma$ are presented in Table 1.
@@
We introduce gradual complexity in terms of covariance matrix choices, starting from known or fixed variance ($M_{1},M_{2}$). For unknown variance, global refers to a common covariance matrix $\Sigma$ for all cluster choices ($M_{3},M_{4},M_{5}$), whereas cluster-specific defines each cluster with a unique covariance matrix $\Sigma_{k}$, along with the mean vectors ($M_{6},M_{7},M_{8}$). The choice of prior distributions maintains conjugacy, while considering hyper-parameters that establish weakly informative priors.
@@ 2.4 Posterior estimation using collapsed VI
We implement a collapsed variational inference (CVI) approach, where we integrate out (or collapse) the stick-breaking parameters $\{V_{k}\}$, following Kurihara et al. (2007). This yields a distribution of the latent allocation vectors $\{z_{n}\}$ depending only on the DP concentration parameter $\alpha$:
@@
(Blei et al., 2017). Although having theoretically “infinite“ prior choices for $k$, truncated variational distributions with an upper limit $k=K$ enable practical implementation (Ishwaran and James, 2001).
For posterior estimation, we use exponential families of distributions for $q(\cdot)$ and apply coordinate ascent algorithm to update the variational hyper-parameters (Bishop and Nasrabadi, 2006). It ensures guarantee of convergence and provides a closed form update $q^{*}(\cdot)$ that maximizes the Evidence lower bound (ELBO) (or equivalently minimizing the information-theoretic Kullbach-Leibler (KL) divergence)
@@
where $\theta_{j}\in\mathbf{\Theta}=\{\alpha,\{\mu_{k}\},\{\Sigma_{k}\},\{z_{n}\}\}$ and $\theta^{-j}=\mathbf{\Theta}\backslash\theta_{j}$ represents the set of all except the $j^{th}$ parameter (Blei and Jordan, 2006; Bishop and Nasrabadi, 2006). Appropriate prior as well as variational distribution choices facilitate the update of $q^{*}(\cdot)$ hyper-parameters. These updates are used to perform posterior inference for the unknown $\{z_{n}\},\{\theta_{i}\}$, thus estimating data cluster as well as distributional partitions corresponding to the cluster allocations.